# Language Supervised Human Action Recognition with Salient Fusion: Construction Worker Action Recognition as a Use-Case

### v1

@@ Abstract
Detecting human actions is a crucial task for autonomous robots and vehicles, often requiring the integration of various data modalities for improved accuracy. In this study, we introduce a novel approach to Human Action Recognition (HAR) based on skeleton and visual cues. Our method leverages a language model to guide the feature extraction process in the skeleton encoder. Specifically, we employ learnable prompts for the language model conditioned on the skeleton modality to optimize feature representation. Furthermore, we propose a fusion mechanism that combines dual-modality features using a salient fusion module, incorporating attention and transformer mechanisms to address the modalities’ high dimensionality. This fusion process prioritizes informative video frames and body joints, enhancing the recognition accuracy of human actions. Additionally, we introduce a new dataset tailored for real-world robotic applications in construction sites, featuring visual, skeleton, and depth data modalities, named VolvoConstAct. This dataset serves to facilitate the training and evaluation of machine learning models to instruct autonomous construction machines for performing necessary tasks in the real world construction zones. To evaluate our approach, we conduct experiments on our dataset as well as three widely used public datasets, NTU-RGB+D, NTU+RGB+D120 and NW-UCLA. Results reveal that our proposed method achieves promising performance across all datasets, demonstrating its robustness and potential for various applications. The codes and dataset are available at:
@@ I INTRODUCTION
Autonomous machines and robots, a rapidly growing industry, must interact effectively with humans using visual cues, auditory commands, or manual controllers. For hazardous environments like construction sites, visual instruction signals are particularly effective [1]. Researchers have explored various methods for recognizing human actions, including body pose [2, 3, 4], visual cues [5, 6, 7], or a combination of both [8, 9, 10] using available public datasets [11, 12]. Skeletons provide detailed limb movement information, while videos offer rich contextual cues such as objects. However, skeleton-based methods require additional deep learning techniques and struggle with actions involving object interactions, while video-based methods are susceptible to environmental disturbances like image blur. Combining these modalities addresses each other’s deficiencies, enhancing overall performance.
@@
Figure 1: A worker providing instructions to the Volvo autonomous wheel loader, demonstrating the integration of manual guidance and automated processes for higher operational efficiency and safety.
@@
Similarly, ActionClip [5] applies this principle to image and text pairs, using a temporal transformer architecture [18] integrated with CLIP to enhance HAR. In both GAP and ActionClip frameworks, the language model acts as a knowledge engine, aiding training and enriching the dataset’s contextual understanding. However, effectively combining these two modalities under the supervision of language models can lead to significant improvements in the field of HAR. In this paper, we introduce a novel end-to-end methodology that integrates human skeleton data and video for HAR, supervised by a language model for global skeleton and limb-based features. Our studies show that our visual modules performs well independently, whereas the skeleton modality relies more on text supervision. The CoOp framework [19] highlights the impact of prompt selection on model performance, advocating for learnable prompt vectors combined with action labels instead of static prompts. CoCoOp [20] further refines this by conditioning the learnable prompt on the other modality. We adopt this strategy to enhance supervision on our skeleton encoder using dynamically adjusted prompts during training.
@@
Integrating features from diverse modalities is crucial for enhancing model accuracy. Traditionally, this has been done using summation [21] or CNNs [22]. However, recent advances show that transformers are highly effective for this purpose [8, 9]. We propose a novel transformer-based fusion approach that prioritizes key features from different modalities. Our method incorporates a fine-grained down-sampling module to highlight valuable data segments and utilizes attention mechanisms to enhance their importance, leading to improved HAR performance.
@@
In addition, our method for Human Activity Recognition (HAR) in construction sites addresses the growing interest in this area, driven by its potential to enhance productivity through the analysis of operations [23, 24] and to improve safety, such as by predicting worker fatigue [25]. A significant challenge in this field is the limited availability of multi-modal datasets tailored for machine instruction in construction environments. Previous research has relied on datasets that are not specifically adapted to these settings. To overcome this limitation, we propose a new dataset specifically designed for HAR in construction workers. This dataset aims to improve the robustness and generalization of robotic systems in real-world construction scenarios.
@@
Contributions: (i) We introduce a novel approach to dual-modality HAR, leveraging language models as knowledge engines. (ii) Employing learnable prompts conditioned on the skeleton features, we supervise and enhance global skeleton and limbs feature extraction. (iii) Our innovative fusion technique implicitly identifies the frames and body joints most crucial for HAR, enhancing feature fusion. (iv) We present, for the first time, a dataset tailored for construction sites to recognize human actions for instructing machines. This dataset aims to enhance safety and efficiency on construction sites by providing data to monitor worker actions, predict hazards, and optimize workflows.
@@ II-A Machine Learning for HAR
Human motion studies and HAR can generally be categorized into skeleton-based, video-based, and skeleton-video-based methods. Skeleton-based methods primarily use CNNs [26], graph-based approaches [27, 28], or attention mechanisms [29, 30, 31] to model joint relationships and skeleton motion. CNN-based methods [4] employ 3D heat maps generated from 2D skeletons instead of GCNs to represent human body structure, followed by CNN layers for further processing. This approach demands significant memory, as each body joint requires a separate image layer. Video-based methods traditionally utilize CNNs as effective local feature learners [32] or combine RGB images with optical flow streams [33]. Notable methods in this domain include Action Machine [34], $\pi$-ViT [35], and DVANet [36]. Skeleton-Visual-based methods leverage multiple modalities to enhance HAR performance [37, 38, 21, 8, 9]. MMNet [21] integrates skeleton joints and bones with RGB images as spatio-temporal regions of interest, enhancing HAR by focusing on key image regions linked to specific joints, though requiring extra pre-processing. Also, STAR-Transformer [8] and STAR++ [9] use spatio-temporal cross-attention to integrate RGB images and skeleton data.
@@ III Methodology
In this section, we explain the technical details of our proposed model. As depicted in Fig. 2, We introduce a skeleton-video based HAR model, which integrates language models as a knowledge engine to supervise the skeleton encoder and the incorporation of learnable prompts (LP) facilitates enhanced utilization of the language models. Fig. 2(a) shows the main model structure. In the training phase, our model is fed with a sequence of RGB images ($V\in{\mathbb{R}}^{T\times H\times W\times 3}$) capturing human actions, alongside the relative sequence of human 3D skeleton joints ($J\in{\mathbb{R}}^{T\times j\times N\times 3}$). Here, $T$ represents number of frames, $j$ the number of skeleton joints and $N$ the number of humans. Also, $H$ and $W$ represent image height and width. Two separate encoders are employed to process data from the visual and skeleton modalities. Our GC-MTC-based skeleton encoder is responsible for learning the skeleton modality. Also, the visual encoder includes a pre-trained ViT-based image encoder and a transformer-based temporal module to learn the relation between the image features. To further guide the learning process, a pre-trained transformer-based text encoder, namely BERT[51] or CLIP-text-encoder [17] is employed to supervise both the global skeleton features and those specific to individual limbs. This is achieved through the use of separate learnable prompt vectors, one for the overall skeletal structure, and one for each distinct body limb section (head, hands, hip, and legs). To enhance the effectiveness of the text encoder’s guidance, the learnable prompts are conditioned on the extracted skeleton features. This is accomplished by adding a feature vector derived from the skeleton data using the Meta-Net module with each learnable prompt vector.
@@
where $y$ is a one-hot vector of the ground-truth labels and $p_{\theta}(x)$ is model’s predicted probability distribution based on the fused features ($x$). Also, in order to calculate the $\mathcal{L}_{\text{cont}}$, it is necessary to first calculate the contrasting text-skeleton vector pairs for each picked batch of data in two directions.
@@
where $s$ and $t$ represent extracted skeleton and text features. Also, sim function is a cosine similarity matrix defined in between the two features and $\tau$ denotes the temperature matrix. The $\mathcal{L}_{\text{cont}}$ is then calculated as following:
@@ III-C Text Encoder
Learnable Prompts. To optimize text encoder performance, we use learnable vectors for prompts rather than pre-defined ones. As shown in Fig. 2(a), the prompts are formatted as $V_{1}V_{2}...[CLASS]...V_{M}$, where each $V_{M}$ (with $m\in{1,\ldots,M}$) is a 512-dimensional vector (for CLIP), initially set to ”X”. The $[CLASS]$ label can appear at the beginning, middle, or end of the prompt, and $M$ represents the number of learnable tokens. To enhance supervision of skeleton features, we use distinct learnable prompts for each body section. A Meta-Net encoder, $h_{\theta}(.)$, parameterized by $\theta$, incorporates a vector derived from the skeleton features into these prompts by adding them together. The Meta-Net consists of a linear layer, a ReLU activation, and another linear layer. Each context token is then obtained by $V_{M}(s)=V_{M}+\pi$, where $s$ denotes skeleton features and $\pi=h_{\theta}(s)$. This approach improves guidance for the text encoder.
@@ V-A Datasets
VolvoConstAct dataset (ours) comprises RGB images, depth data, and 2D and 3D human skeletons. These skeletons are extracted using the human detection capabilities of the ZED 2i camera. The dataset features 17 distinct actions performed by construction workers instructing a vehicle. Motions are captured in 20 unique videos containing specific details and each recorded motion is divided into 5 sections. We use 75% of the total recorded videos from each action as the training set and the remaining 25% as the test set. Additional details about the dataset, is provided in Table I.
@@
Camera Model Stereo RGB+D ZED 2i Camera
@@
NTU-RGB+D [54] is a public dataset widely used for video and/or skeleton based human motion studies and contains 56,880 skeletal action sequences for 60 action classes. The benchmarks for evaluation include Cross-Subject (XSub) and Cross-View (XView). In the XSub evaluation, both training and testing sets comprise data sourced from 20 distinct subjects, ensuring generalizability of the models trained on this dataset. Conversely, the XView evaluation leverages samples captured by cameras 2 and 3, totaling 37,920 sequences, for training purposes. The remaining 18,860 sequences, captured by camera 1, are reserved for the test set in this evaluation framework.
NTU-RGB+D 120 [11] extends the NTU-RGB+D dataset with an additional 57,367 video and skeleton sequence captured for an extra 60 action classes creating 120 action classes in total. Two benchmark evaluations exist for this dataset including Cross-Subject (XSub) and Cross-Setup (XSet) settings.
@@ V-B Implementation Details
In our experiments, using entire images as the visual modality did not yield optimal performance. However, applying a cropping technique to focus solely on the human subject significantly improved model performance and reduced the data volume. This method is particularly effective with ViT as the visual encoder, enabling the attention mechanism to focus on relevant objects related to human actions, such as a glass for the “drinking water” action. We used YOLOv5 [48] for human detection and cropping. For actions involving multiple subjects, all detected humans and the intervening area were included. During training and testing, all image frames were resized to uniform dimensions. For skeleton data from the NTU-RGB+D and NTU-RGB+D120 datasets, we applied the denoising techniques described in [55, 3]. For the NW-UCLA dataset, we followed the standard procedure outlined in [16]. Also, we used 16 frames for both RGB images and skeleton modalities, uniformly selecting data from the total frames of the video and the associated skeleton frames, as detailed in [8, 9] and [56].
@@ V-C Results
In Table II, we present a comparative analysis of our method against several state-of-the-art (SOTA) approaches on the NTU-RGB+D, NTU-RGB+D120 and NW-UCLA datasets. For a fair comparison, we primarily focus and restrict our observation in Table II on methods that do not require pre-training of their modality encoders. As depicted in Table II, we achieve better or comparable performance with respect to the SOTA methods. Specifically, for NTU-RGB+D, we improve model performance by 1% compared to the closest SOTA method [9] for the NTU-60 XSub benchmark. For the remaining benchmarks, our results are comparable in terms of accuracy percentage. It is important to note that we avoid fine-tuning our models for the other benchmarks, which accounts for the slight difference between our results and the best SOTA results.
Additionally, Fig. 3 presents a qualitative result showing selected images and skeleton joints for a sequence depicting a human performing a “waving” action. The most effective visual features are associated with frames exhibiting significant hand movements, while the skeleton features related to the waving arm have the greatest impact on action detection.
@@
Ours R+P (✗) 93.7 97.3 90.3 89.4 95.8

### v2

# LS-HAR: Language Supervised Human Action Recognition with Salient Fusion, Construction Sites as a Use-Case

@@ Abstract
Detecting human actions is a crucial task for autonomous robots and vehicles, often requiring the integration of various data modalities for improved accuracy. In this study, we introduce a novel approach to Human Action Recognition (HAR) using language supervision named LS-HAR based on skeleton and visual cues. Our method leverages a language model to guide the feature extraction process in the skeleton encoder. Specifically, we employ learnable prompts for the language model conditioned on the skeleton modality to optimize feature representation. Furthermore, we propose a fusion mechanism that combines dual-modality features using a salient fusion module, incorporating attention and transformer mechanisms to address the modalities’ high dimensionality. This fusion process prioritizes informative video frames and body joints, enhancing the recognition accuracy of human actions. Additionally, we introduce a new dataset tailored for real-world robotic applications in construction sites, featuring visual, skeleton, and depth data modalities, named VolvoConstAct. This dataset serves to facilitate the training and evaluation of machine learning models to instruct autonomous construction machines for performing necessary tasks in real-world construction sites. To evaluate our approach, we conduct experiments on our dataset as well as three widely used public datasets: NTU-RGB+D, NTU-RGB+D 120, and NW-UCLA. Results reveal that our proposed method achieves promising performance across all datasets, demonstrating its robustness and potential for various applications. The code, dataset, and demonstration of real-machine experiments are available at:
@@ I INTRODUCTION
Autonomous machines and robots, a rapidly growing industry, must interact effectively with humans using visual cues, auditory commands, or manual controllers. For hazardous environments like construction sites, visual instruction signals are particularly effective [1]. Researchers have explored various methods for recognizing human actions, including body pose [2, 3, 4], visual cues [5, 6, 7], or a combination of both [8, 9] using available public datasets [10, 11]. Skeletons provide detailed limb movement information, while videos offer rich contextual cues such as objects. However, skeleton-based methods require additional deep learning techniques and struggle with actions involving object interactions, while video-based methods are susceptible to environmental disturbances like image blur. Combining these modalities addresses each other’s deficiencies, enhancing overall performance.
@@
Figure 1: A construction worker providing instructions to the Volvo autonomous wheel loader, demonstrating the integration of manual guidance and automated processes for higher operational efficiency and safety.
@@
Similarly, ActionClip [5] applies this principle to image and text pairs, using a temporal transformer architecture [17] integrated with CLIP to enhance HAR. In both GAP and ActionClip frameworks, the language model acts as a knowledge engine, aiding training and enriching the dataset’s contextual understanding. However, effectively combining these two modalities under the supervision of language models can lead to significant improvements in the field of HAR. In this paper, we introduce a novel end-to-end methodology that integrates human skeleton data and video for HAR, supervised by a language model for global skeleton and limb-based features. Our studies show that our visual modules perform well independently, whereas the skeleton modality relies more on text supervision. The CoOp framework [18] highlights the impact of prompt selection on model performance, advocating for learnable prompt vectors combined with action labels instead of static prompts. CoCoOp [19] further refines this by conditioning the learnable prompt on the other modality. We adopt this strategy to enhance supervision on our skeleton encoder using dynamically adjusted prompts during training.
@@
Integrating features from diverse modalities is crucial for enhancing model accuracy. Traditionally, this has been done using summation [20] or CNNs [21]. However, recent advances show that transformers are highly effective for this purpose [8]. We propose a novel transformer-based fusion approach that prioritizes key features from different modalities. Our method incorporates a fine-grained down-sampling module to highlight valuable data segments and utilizes attention mechanisms to enhance their importance, leading to improved HAR performance.
@@
In addition, our method for Human Activity Recognition (HAR) in construction sites addresses the growing interest in this area, driven by its potential to enhance productivity through the analysis of operations [22, 23] and to improve safety, such as by predicting worker fatigue [24]. A significant challenge in this field is the limited availability of multi-modal datasets tailored for machine instruction in construction environments. Previous research has relied on datasets that are not specifically adapted to these settings. To overcome this limitation, we propose VolvoConstAct, a new dataset designed for HAR in construction workers. This dataset aims to improve the robustness and generalization of robotic systems in real-world construction environments.
@@
Contributions: (i) We introduce a novel approach to dual-modality HAR, leveraging language models as knowledge engines. (ii) Employing learnable prompts conditioned on the skeleton features, we supervise and enhance global skeleton and limb feature extraction. (iii) Our innovative fusion technique implicitly identifies the frames and body joints most crucial for HAR, enhancing feature fusion. (iv) We present, for the first time, a dataset tailored for construction sites to recognize human actions for instructing machines. This dataset aims to enhance safety and efficiency on construction sites by providing data to monitor worker actions, predict hazards, and optimize workflows.
@@ II-A Machine Learning for HAR
Human motion studies and HAR can generally be categorized into skeleton-based, video-based, and skeleton-video-based methods. Skeleton-based methods primarily use CNNs [25], graph-based approaches [26, 27, 28], or attention mechanisms [29, 30, 31] to model joint relationships and skeleton motion. CNN-based methods [4] employ 3D heat maps generated from 2D skeletons instead of GCNs to represent human body structure, followed by CNN layers for further processing. This approach demands significant memory, as each body joint requires a separate image layer. Video-based methods traditionally utilize CNNs as effective local feature learners [32] or combine RGB images with optical flow streams [33]. Notable methods in this domain include Action Machine [34], $\pi$-ViT [35], and DVANet [36]. Skeleton-Visual-based methods leverage multiple modalities to enhance HAR performance [37, 38, 20, 8]. MMNet [20] integrates skeleton joints and bones with RGB images as spatio-temporal regions of interest, enhancing HAR by focusing on key image regions linked to specific joints, though it requires additional pre-processing. Also, STAR-Transformer [8] uses spatio-temporal cross-attention to integrate RGB images and skeleton data.
@@ II-B Applications of Language Models in HAR
On a related note, Xu et al. [39] leveraged an LLM assistant to train a skeleton-based model. It maps LLM knowledge into a priori global relationship (GPR) and a priori category relationship (CPR) topologies. GPR refines bone representations to highlight key node information, while CPR simulates category priors in the brain, using a PC-AC module for additional supervision to enhance class-distinguishable features.
@@ III Methodology
In this section, we explain the technical details of our proposed model. As depicted in Fig. 2, we introduce a skeleton-video based HAR model, which integrates language models as a knowledge engine to supervise the skeleton encoder and the incorporation of learnable prompts (LP) facilitates enhanced utilization of the language models. Fig. 2(a) shows the main model structure. In the training phase, our model is fed with a sequence of RGB images ($V\in{\mathbb{R}}^{T\times H\times W\times 3}$) capturing human actions, alongside the relative sequence of human 3D skeleton joints ($J\in{\mathbb{R}}^{T\times j\times N\times 3}$). Here, $T$ represents the number of frames, $j$ the number of skeleton joints, and $N$ the number of humans. Also, $H$ and $W$ represent image height and width. Two separate encoders are employed to process data from the visual and skeleton modalities. Our GC-MTC-based skeleton encoder is responsible for learning the skeleton modality. Also, the visual encoder includes a pre-trained ViT-based image encoder and a transformer-based temporal module to learn the relation between the image features.
To further guide the learning process, a pre-trained transformer-based text encoder, such as BERT [52] or CLIP-text-encoder [16], is employed to supervise both global skeleton features and those specific to individual limbs. This is achieved using separate learnable prompt vectors: one for the overall skeletal structure and one for each distinct body limb section (head, hands, hips, and legs). To enhance the effectiveness of the text encoder’s guidance, the learnable prompts are conditioned on the extracted skeleton features. This is accomplished by adding a feature vector derived from the skeleton data using the Meta-Net module with each learnable prompt vector.
@@
where $y$ is a one-hot vector of the ground-truth labels and $p_{\theta}(x)$ is the model’s predicted probability distribution based on the fused features ($x$). Also, in order to calculate the $\mathcal{L}_{\text{cont}}$, it is necessary to first calculate the contrasting text-skeleton vector pairs for each picked batch of data in two directions.
@@
where $s$ and $t$ represent extracted skeleton and text features. Also, sim function is a cosine similarity matrix defined between the two features, and $\tau$ denotes the temperature matrix. The $\mathcal{L}_{\text{cont}}$ is then calculated as follows:
@@ III-C Text Encoder
Learnable Prompts. To optimize text encoder performance, we use learnable vectors for prompts rather than pre-defined ones. As shown in Fig. 2(a), the prompts are formatted as $V_{1}V_{2}...[CLASS]...V_{M}$, where each $V_{M}$ (with $m\in{1,\ldots,M}$) is a 512-dimensional vector (for CLIP), initially set to “X”. The $[CLASS]$ label can appear at the beginning, middle, or end of the prompt, and $M$ represents the number of learnable tokens. To enhance supervision of skeleton features, we use distinct learnable prompts for each body section. A Meta-Net encoder, $h_{\theta}(.)$, parameterized by $\theta$, incorporates a vector derived from the skeleton features into these prompts by adding them together. The Meta-Net consists of a linear layer, a ReLU activation, and another linear layer. Each context token is then obtained by $V_{M}(s)=V_{M}+\pi$, where $s$ denotes skeleton features and $\pi=h_{\theta}(s)$. This approach improves guidance for the text encoder.
@@ V-A Datasets
VolvoConstAct dataset (ours) comprises RGB images, depth data, and 2D and 3D human skeletons. These skeletons are extracted using the human detection capabilities of the ZED 2i camera. The dataset features 17 distinct actions performed by construction workers instructing a vehicle. Motions are captured in 20 unique videos containing specific details, and every recorded motion is divided into 5 sections. We use 75% of the total recorded videos from each action as the training set and the remaining 25% as the test set. Additional details about the dataset are provided in Table I.
@@
Camera Model Stereo RGB+D ZED 2i Camera  [^1] https://www.stereolabs.com/en-se/store/products/zed-2i
@@
NTU-RGB+D [55] is a public dataset widely used for video and/or skeleton based human motion studies and contains 56,880 skeletal action sequences for 60 action classes. The benchmarks for evaluation include Cross-Subject (XSub) and Cross-View (XView). In the XSub evaluation, both training and testing sets comprise data sourced from 20 distinct subjects, ensuring generalizability of the models trained on this dataset. Conversely, the XView evaluation leverages samples captured by cameras 2 and 3, totaling 37,920 sequences for training purposes. The remaining 18,860 sequences, captured by camera 1, are reserved for the test set in this evaluation framework.
NTU-RGB+D 120 [10] extends the NTU-RGB+D dataset with an additional 57,367 video and skeleton sequences, capturing an extra 60 action classes, resulting in a total of 120 action classes. Two benchmark evaluations exist for this dataset, including Cross-Subject (XSub) and Cross-Setup (XSet) settings.
@@ V-B Implementation Details
In our experiments, using entire images as the visual modality did not yield optimal performance. However, applying a cropping technique to focus solely on the human subject significantly improved model performance and reduced the data volume. This method is particularly effective with ViT as the visual encoder, enabling the attention mechanism to focus on relevant objects related to human actions, such as a glass for the “drinking water” action. We used YOLOv5 [49] for human detection and cropping. For actions involving multiple subjects, all detected humans and the intervening area were included. During training and testing, all image frames were resized to uniform dimensions. For skeleton data from the NTU-RGB+D and NTU-RGB+D120 datasets, we applied the denoising techniques described in [56, 3]. For the NW-UCLA dataset, we followed the standard procedure outlined in [15]. Also, we used 16 frames for both RGB images and skeleton modalities, uniformly selecting data from the total frames of the video and the associated skeleton frames, as detailed in [8] and [57].
@@ V-C Results
In Table II, we present a comparative analysis of our method against several state-of-the-art (SOTA) approaches on the NTU-RGB+D, NTU-RGB+D120 and NW-UCLA datasets. For a fair comparison, we primarily focus and restrict our observation in Table II on methods that do not require pre-training of their modality encoders. As depicted in Table II, we achieve better or comparable performance with respect to the SOTA methods. Specifically, for NTU60-XSub, we improve model performance by 1% compared to the closest similar SOTA method [9] and 0.7% compared to the best pose-based method [3]. Also, for the NTU60-XView, NTU120-XSub and UCLA benchmarks, our results and the best SOTA are comparable.
It is important to highlight that we fine-tuned our model and hyperparameters specifically for the NTU60-XSub benchmark, which explains the small performance gap between our results and the best SOTA results on other benchmarks. The performance gap between our method and the best SOTA is more pronounced on the NTU120-XSet benchmark. This discrepancy can be attributed to differences in camera setups and view angles between our training and testing datasets. Our approach involves cropping the human from each image frame, which can potentially reduce the similarity between the training and testing samples, thereby impacting model performance. Therefore, this does not indicate a lower overall quality of our presented method compared to the SOTA.
Additionally, methods such as TSMF [9], STAR [8], and STAR++ [58] use heatmaps to learn 2D skeletons, requiring significantly more GPU memory during inference than our skeleton encoder. In contrast, our approach employs GCNs to efficiently process skeleton data, making it more suitable for embedded systems in construction machinery, where GPU memory is often constrained in edge computing devices.
We also trained our model using the VolvoConstAct dataset and compared its performance with some of the SOTA methods that have publicly available open-source code suitable for embedded systems in construction machines. We achieved better results compared to the SOTA methods as shown in Table III. It is important to mention that our goal is to establish a baseline for action recognition of construction workers using our dataset. Enhanced performance for the skeleton modality on VolvoConstAct could be achieved with additional denoising and pre-processing, as we used raw skeleton data to test the model’s robustness and establish our evaluations as a baseline. Additionally, our high performance on video only accuracy suggests that future studies could benefit from focusing solely on the visual modality, which is more advantageous for long-range HAR.
Also, Fig. 3 presents a qualitative result showing selected images and skeleton joints for a sequence depicting a human performing a “waving” action. The most effective visual features are associated with frames exhibiting significant hand movements, while the skeleton features related to the waving arm have the greatest impact on action detection.
@@
Ours R+P (✗) 93.7 97.3 90.3 89.4 96.9
TABLE III: Model performance comparison with SOTA on VolvoConstAct (Volvo) dataset.
Method VolvoConstAct (%)
VideoMAE v2 small [62] 70.5
VideoMAE v2 [62] 86.2
VideoSwin [63] 96.0
PoseC3D [4] 94.8
Ours (Video Only) 92.9
Ours with ViT-B/32 91.5
Ours 96.4